OpenStudy (anonymous):

A red kite is 100 feet off the ground and is rising at 8 feet per second. A blue kite is 180 feet off the ground and is rising at 5 feet per second. How long will it take for the red kite to be higher than the blue kite? Round your answer to the nearest second

OpenStudy (anonymous):

we first find when they will be at equal height let them be at same height after x seconds this means 100 + 8x = 180 + 5x 8x - 5x = 180 - 100 3x = 80 x = 80/3 = 26.67 so they will be at same height after 26.67 seconds so it will take 27 seconds for the red kite to be higher than the blue kite

OpenStudy (anonymous):

Red Kite: 100/8=12.5 Blue Kite: 180/36 36-12.5= 23.5 seconds or rounded= 24 seconds.